# Direct Quantized Training of Language Models with Stochastic Rounding

### v2

@@ Abstract
Although recent quantized Large Language Models (LLMs), such as BitNet, have paved the way for significant reduction in memory usage during deployment with binary or ternary weights, training these models still demands substantial memory footprints. This is partly because high-precision (i.e., unquantized) weights required for straight-through estimation must be maintained throughout the whole training process. To address this, we explore directly updating the quantized low-precision weights without relying on straight-through estimation during backpropagation, aiming to save memory usage during training. Specifically, we employ a stochastic rounding technique to minimize the information loss caused by the use of low-bit weights throughout training. Experimental results on our LLaMA-structured models of various sizes indicate that (1) training with only low-precision weights is feasible even when they are constrained to ternary values; (2) extending the bit width to 8 bits achieves performance on par with BitNet b1.58; (3) our models remain robust to precision scaling and memory reduction, showing minimal performance degradation when moving from FP32 to lower-memory environments (BF16/FP8); and (4) our models also support inference using ternary weights, showcasing their flexibility in deployment.  [^1] Work in progress. Code is available at https://github.com/KYuuto1006/DQT
Direct Quantized Training of Language Models with Stochastic Rounding
Kaiyan Zhao1, Tsuguchika Tabaru2, Kenichi Kobayashi2, Takumi Honda2, Masafumi Yamazaki2, Yoshimasa Tsuruoka1 1The University of Tokyo, Tokyo, Japan 2Fujitsu Limited, Kawasaki, Japan 1{zhaokaiyan1006, yoshimasa-tsuruoka}@g.ecc.u-tokyo.ac.jp 2{tabaru, kenichi, honda.takumi, m.yamazaki}@fujitsu.com
@@ 1 Introduction
Large Language Models (LLMs) have become a promising solution for a wide range of Natural Language Processing (NLP) tasks, including machine translation (Xu et al., 2024a; Wu et al., 2024), summarization (Zhang et al., 2024a), reasoning (Wei et al., 2024; Kojima et al., 2024; Miao et al., 2024) and question answering (OpenAI, 2022, 2023). However, their development is challenged by the need for vast datasets and substantial computational resources, especially as the size of current LLMs continually grows larger (Duan et al., 2024).
Quantization, which involves converting high-precision parameter matrices into lower-precision formats, has emerged as an effective approach for enabling resource-efficient LLMs (Egashira et al., 2024). Traditional quantization methods can be divided into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). PTQ reduces the bit precision of weight matrices in an already pre-trained LLM (Banner et al., 2019; Frantar et al., 2023), while QAT incorporates quantization during training, enabling the model to adapt to low-bit precision throughout the learning process (Jacob et al., 2018).
Figure 1: Comparison of the training process for BitNet and our modified one. Upper: The training process for BitNet, where the original high precision weights are updated with the straight-through estimator in backward process. Lower: We directly update the low precision weights with stochastic rounding, eliminating the need to quantize the weight matrices in each training step and keeping weight matrices always at low-bit. For clarity, we illustrate our method using ternary examples. We present an example of 8-bit quantization in Appendix, Figure 8.
@@ 2 Related Works
Quantization for deep neural networks has a history spanning nearly ten years, with researchers initially compressing networks to reduce memory usage and computational load while maintaining accuracy (Rastegari et al., 2016; Hubara et al., 2018). Recent quantization methods for LLMs can be divided into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT).
2.1 Post-Training Quantization
@@
2.2 Quantization-Aware Training
To bridge the gap between high-precision parameter training and quantization, QAT incorporates the quantization of model parameters during the training process. One of the initial applications of QAT to LLMs comes from Liu et al. (2024), who propose a data-free distillation-based method and quantize the model to 4 bits. Xu et al. (2024b) further expand distillation-based methods to binary quantization by introducing limited trainable vectors. More recently, BitNet (Wang et al., 2023; Ma et al., 2024) is proposed, achieving training from scratch QAT with weight values constrained to {-1,0,1}. However, due to the non-differentiability of the quantization process, special gradient approximation methods like the straight-through estimator (STE) (Bengio et al., 2013) are commonly employed during training. While effective, this approach often results in slower training processes and increased computational memory overhead since high-precision weights are always maintained during the training process. These inefficiencies become particularly pronounced when scaling QAT to larger models, limiting the practical training of QAT in real-world use cases. To address this, in this work, we explore the potential for more efficient and practical QAT methods that have significantly lower memory dependency compared to traditional QAT approaches.
@@ 3.1 Stochastic Rounding
The idea of stochastic rounding (SR) originates from Von Neumann and Goldstine (1947), and is initially used for reducing bias in numerical computations and recently for deep learning models (Gupta et al., 2015). It is a rounding technique that probabilistically rounds values to the nearest representable precision based on their distance from those values. Given a high precision value $x$, stochastic rounding can be defined as the following equation Gupta et al. (2015); Markov et al. (2023); Zhang et al. (2024b):
@@ 3.2 Modified Training Process
Next, we continue to explain the details of DQT. As shown in Figure 1 (b), we start training from low-precision weight matrices instead of high-precision ones. We achieve this through utilizing AbsMean Quantization following (Ma et al., 2024) on a randomly initiated weight matrix $W$. The absolute mean value for $W$ can be represented as:
@@
where $w_{i}$ is the $i_{th}$ element in $W$. For $n$-bit quantization, $Q_{n}=-2^{n-1}$ and $Q_{p}=2^{n-1}-1$ are then determined to constrain the range of quantization, and the scaling factor $s$ can be defined as:
@@
Finally, the quantized weights $\tilde{W}$ can be expressed in the following equation:
$\tilde{W}=\mathrm{Clip}[\mathrm{Round}(W\cdot s),Q_{n},Q_{p}]/s,$ (4)
where $\mathrm{Clip}()$ function assures that all the values are in the range of $[Q_{n},Q_{p}]$ and $\mathrm{Round}()$ returns the nearest integer. This allows us to constrain the quantized $\tilde{W}$ to $n$-bits (INTn). While for the inputs and activations, we follow the settings introduced in BitNet (Wang et al., 2023; Ma et al., 2024) and quantize them into 8 bits.
After computing the language modeling cross-entropy loss based on $\tilde{W}$ and inputs in each training step, we first allow the optimizer to calculate $W^{\prime}$, the dense weight matrix intended for updating. In traditional QAT, the straight-through estimator is applied, and $W^{\prime}$ replaces the original high-precision $W$, subsequently undergoing the quantization process from Equation (2) to Equation (4) again in the next training step. However, in our approach, we simplify this process by directly applying stochastic rounding on $W^{\prime}$ using Equation (1):
$\widetilde{W}=\mathrm{SR}(W^{\prime}),$ (5)
to make sure it maintains $n$-bits (INTn) without requiring the retention of high-precision weights, thus getting rid of the straight-through estimator and skipping the process from Equation (2) to Equation (4). During our modified DQT, we directly replace $\tilde{W}$ with $\widetilde{W}$ and proceed to the next training step, ensuring that the weight matrices are always constrained to $n$-bits throughout training, which makes the biggest difference from traditional QAT.
@@ 4.1 Implementation Setup
We follow the architecture of LLaMA2 Touvron et al. (2023); Dubey et al. (2024) and initialize models with sizes ranging from 130M to 320M and 1B parameters. Detailed model specifications and configurations are provided in Appendix A.1. For the main experiments, we use AdamW Loshchilov and Hutter (2019) as the optimizer. For DQT models, we apply modified versions of the optimizers that incorporate stochastic rounding.
@@
We use two types of datasets to pretrain the models. The first is the English Wikipedia dataset (20231101.en) [^2] https://huggingface.co/datasets/wikimedia/wikipedia. For the 1B models, we further pre-train them using a larger dataset containing 10B tokens from the FineWeb dataset (Penedo et al., 2024) [^3] https://huggingface.co/datasets/HuggingFaceFW/fineweb/viewer/sample-10BT. We split 1% of the data as the corresponding development set. Models are trained for one epoch with a cosine scheduler applied and a 2000 step warm-up. Refer to Appendix A.1 for details.
@@
Besides reporting the training loss and perplexity results, we perform zero-shot evaluation on general language modeling tasks using the lm_eval benchmark Gao et al. (2024). Specifically, we evaluate on WinoGrande Sakaguchi et al. (2019), ARC Clark et al. (2018), PIQA Bisk et al. (2020) and SciQ Welbl et al. (2017), which span a range of reasoning and question answering tasks across diverse domains.
Figure 2: Comparison of our DQT and other baselines accross different model sizes and training datasets. The horizontal axis represents the training steps while the vertical axis represents the training loss. As model size increases, the performance of our DQT models, especially DQT 8 bit, become more comparable to and even better than the reproduced BitNet b1.58.
@@
We pre-train the 130M models on 4 NVIDIA A100 80GB GPUs, the 330M models on 8 NVIDIA GH200 Grace Hopper Superchips, and the 1B models on 16 GH200 Superchips. To remain hardware-agnostic, we simulate our quantization approach under FP32, BF16, and FP8 environments. We use MS-AMP [^4] https://github.com/Azure/MS-AMP for FP8 experiments with optimization level set to MS-AMP O2. While we recognize that FP8 is still not true ternary precision, it offers a practical compromise under current hardware constraints, as done in many prior works. Refer to Appendix A.1 for more details.
@@ 4.2 Main Results
We first present the training loss for DQT variants, our reproduced LLaMA (FP32) and BitNet b1.58 in Figure 2, including different model sizes and training datasets under FP32. In all subfigures of Figure 2, the blue line represents the ternary implementation of DQT, where weight matrices are always constrained to {-1, 0, 1}. The green line denotes the 8 bit (INT8) implementation of DQT. The orange line corresponds to the reproduced BitNet b1.58, which utilizes high-precision information during training. Finally, the red line indicates the standard FP32 LLaMA implementation.
We first examine the blue lines: although there remains a performance gap compared to higher-precision models, the ternary DQT implementation still demonstrates the ability to converge. Across all model sizes and training datasets, we observe that standard FP32 models (red lines) consistently achieve the best performance, as they are not subject to any quantization. BitNet (orange lines) consistently delivers performance close to FP32, benefiting from high-precision updates during training. Finally, for our DQT 8 bit implementation (green lines), the performance gap with BitNet b1.58 narrows as model size increases. Particularly, in the 1B models, DQT 8 bit models surpass BitNet b1.58, and this trend persists even when training on the larger FineWeb dataset. A clearer non-logarithmic comparison between the 1B DQT 8-bit model and BitNet b1.58 is provided in Appendix, Figure 10.
@@ 4.3 Low Memory Experiments
We additionally conduct two types of experiments to evaluate DQT’s performance given reduced GPU memory, simulating resource-limited environments. First, we assess performance in low-precision formats using BF16 and FP8. Second, we examine the effect of memory-efficient optimizers, as standard AdamW maintains two high-precision states (momentum and variance) for each parameter, contributing significantly to memory usage. To demonstrate DQT’s low reliance on high-precision information, we specifically choose Adafactor (Shazeer and Stern, 2018), which eliminates the need for fully storing these additional states. Note that we conduct experiments with Adafactor in BF16 and FP8 formats, as its benefits are most pronounced in low-precision environments.
Figure 3 presents the GPU memory usage and corresponding development set loss at the end of training for 130M and 1B models trained on the Wikipedia dataset. The y-axis indicates the percentage of GPU memory used on a single GH200 GPU under our experimental settings. We explain how we acquire the percentage in Appendix A.3. As shown in the figure, BitNet models experience clear performance degradation as GPU memory usage decreases in BF16 and FP8 settings, for both 130M and 1B scales. This is expected, since BitNet still relies on high-precision information for effective weight updates. In contrast, our DQT 8-bit models maintain robust performance under reduced precision. For both 130M and 1B models, the performance drop is less than 0.1 in loss, demonstrating DQT’s resilience to lower GPU memory usage and its minimal dependence on high-precision information. This robustness persists even when applying memory-efficient optimizers such as Adafactor in both BF16 and FP8 settings. These results suggest that DQT models can be effectively trained even in memory-constrained environments, showing consistent performance across both low-precision formats and lightweight optimization methods. We believe that DQT can be further extended to even lower-precision settings while maintaining its effectiveness.
@@ 4.4 Impact of the Bit Width in DQT
Figure 4: Comparison of different bit widths in DQT. Higher $n$-bit results in better performance.
Next, we proceed to examine the impact of varying the number of bits in DQT. Specifically, we experiment with $n$ in {1.58, 3, 4, 8} on 130M models trained with the Wikipedia dataset and 1B models trained with the FineWeb dataset. The corresponding results are presented in Figure 4. We select the optimal learning rate for each model.

### v3

@@ Abstract
Although recent quantized Large Language Models, such as BitNet, have paved the way for significant reduction in memory usage during deployment with binary or ternary weights, training these models still demands substantial memory footprints. This is partly because high-precision (i.e., unquantized) weights required for straight-through estimation must be maintained throughout the whole training process. To address this, we explore directly updating the quantized low-precision weights without relying on straight-through estimation during backpropagation, aiming to save memory usage during training. Specifically, we employ a stochastic rounding technique to minimize the information loss caused by the use of low-bit weights throughout training. Experimental results on our LLaMA-structured models of various sizes indicate that (1) training with only low-precision weights is feasible even when they are constrained to ternary values; (2) extending the bit width to 8 bits achieves performance on par with BitNet b1.58; (3) our models remain robust to precision scaling and memory reduction, showing minimal performance degradation when moving from FP32 to lower-memory environments (BF16/FP8); and (4) our models also support inference using ternary weights, showcasing their flexibility in deployment. [^1] Code is available at https://github.com/KYuuto1006/DQT.
keywords:
Large Language Models; Quantization-Aware Training.
editors: Hung-yi Lee and Tongliang Liu
@@ 1 Introduction
Large Language Models (LLMs) have become a promising solution for a wide range of Natural Language Processing (NLP) tasks, including machine translation (Xu et al., 2024a; Wu et al., 2024; Miao et al., 2025), reasoning (OpenAI, 2022, 2024; Kojima et al., 2024; Miao et al., 2024) and multimodal understanding (Huang et al., 2025; Mao et al., 2025). However, their development is challenged by the need for vast datasets and substantial computational resources, especially as the size of current LLMs continually grows larger (Duan et al., 2024).
Quantization, which involves converting high-precision parameter matrices into lower-precision formats, has emerged as an effective approach for enabling resource-efficient LLMs. Traditional quantization methods can be divided into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). PTQ reduces the bit precision of weight matrices in an already pre-trained LLM (Banner et al., 2019; Frantar et al., 2023), while QAT incorporates quantization during training, enabling the model to adapt to low-bit precision throughout the learning process (Jacob et al., 2018).
Figure 1: Comparison of the training process for BitNet and our modified one. Upper: The training process for BitNet, where the original high precision weights are updated with the straight-through estimator in backward process. Lower: We directly update the low precision weights with stochastic rounding, eliminating the need to quantize the weight matrices in each training step and keeping weight matrices always at low-bit. We provide an 8-bit example in Supplementary Material, Figure 3.
@@ 2 Related Works
Efficient learning methods for LLMs have become a critical area of research (Zhao et al., 2025). Quantization for deep neural networks has a history spanning nearly ten years, with researchers initially compressing networks to reduce memory usage and computational load while maintaining accuracy (Rastegari et al., 2016; Hubara et al., 2018). Recent quantization methods for LLMs can be divided into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT).
Post-Training Quantization
@@
Quantization-Aware Training
To bridge the gap between high-precision parameter training and quantization, QAT incorporates the quantization of model parameters during the training process. One of the initial applications of QAT to LLMs comes from Liu et al. (2024), who propose a data-free distillation-based method and quantize the model to 4 bits. Xu et al. (2024b) further expand distillation-based methods to binary quantization by introducing limited trainable vectors. More recently, BitNet (Wang et al., 2023; Ma et al., 2024) is proposed, achieving training from scratch QAT with weight values constrained to $\{-1,0,1\}$. However, due to the non-differentiability of the quantization process, special gradient approximation methods like the straight-through estimator (STE) (Bengio et al., 2013) are commonly employed during training. While effective, this approach often results in slower training processes and increased computational memory overhead since high-precision weights are always maintained during the training process. These inefficiencies become particularly pronounced when scaling QAT to larger models, limiting the practical training of QAT in real-world use cases. To address this, in this work, we explore the potential for more efficient and practical QAT methods that have significantly lower memory dependency compared to traditional QAT approaches.
@@ 3.1 Stochastic Rounding
The idea of stochastic rounding (SR) originates from Von Neumann and Goldstine (1947), and is initially used for reducing bias in numerical computations and recently for deep learning models (Gupta et al., 2015). It is a rounding technique that probabilistically rounds values to the nearest representable precision based on their distance from those values. Given a high precision value $x$, stochastic rounding can be defined as the following equation (Gupta et al., 2015; Markov et al., 2023; Zhang et al., 2024):
@@ 3.2 Modified Training Process
Next, we continue to explain the details of DQT. As shown in Figure 1 (b), we start training from low-precision weight matrices instead of high-precision ones. We achieve this through utilizing AbsMean Quantization following Ma et al. (2024) on a randomly initiated weight matrix $W$. The absolute mean value for $W$ can be represented as
@@
where $w_{i}$ is the $i_{th}$ element in $W$. For $n$-bit quantization, $Q_{n}=-2^{n-1}$ and $Q_{p}=2^{n-1}-1$ are then determined to constrain the range of quantization, and the scaling factor $s$ can be defined as
@@
Finally, the quantized weights $\widetilde{W}$ can be expressed in the following equation:
$\widetilde{W}=\mathrm{Clip}[\mathrm{Round}(W\cdot s),Q_{n},Q_{p}]/s,$ (4)
where the $\mathrm{Clip}()$ function assures that all the values are in the range of $Q_{n},Q_{p}]$ and $\mathrm{Round}()$ returns the nearest integer. This allows us to constrain the quantized $\widetilde{W}$ to $n$-bits (INTn). While for the inputs and activations, we follow the settings introduced in BitNet (Wang et al., 2023; Ma et al., 2024) and quantize them into 8 bits.
After computing the language modeling cross-entropy loss based on $\widetilde{W}$ and inputs in each training step, we first allow the optimizer to calculate $W^{\prime}$, the dense weight matrix intended for updating. In traditional QAT, the straight-through estimator is applied, and $W^{\prime}$ replaces the original high-precision $W$, subsequently undergoing the quantization process from Equation (2) to Equation (4) again in the next training step. However, in our approach, we simplify this process by directly applying stochastic rounding on $W^{\prime}$ using Equation (1):
$\widehat{W}=\mathrm{SR}(W^{\prime}),$ (5)
to make sure it maintains $n$-bits (INTn) without requiring the retention of high-precision weights, thus getting rid of the straight-through estimator and skipping the process from Equation (2) to Equation (4). During our modified DQT, we directly replace $\widetilde{W}$ with $\widehat{W}$ and proceed to the next training step, ensuring that the weight matrices are always constrained to $n$-bits throughout training, which makes the biggest difference from traditional QAT.
@@ 4.1 Implementation Setup
We follow the architecture of LLaMA2 (Touvron et al., 2023) and initialize models with sizes ranging from 130M to 320M and 1B parameters. Detailed model specifications and configurations are provided in the Supplementary Material, Section 1.1. For the main experiments, we use AdamW (Loshchilov and Hutter, 2019) as the optimizer. For DQT models, we apply modified versions of the optimizers that incorporate stochastic rounding.
@@
We use two types of datasets to pretrain the models. The first is the English Wikipedia dataset (20231101.en) [^2] https://huggingface.co/datasets/wikimedia/wikipedia. For the 1B models, we further pre-train them using a larger dataset containing 10B tokens from the FineWeb dataset (Penedo et al., 2024) [^3] https://huggingface.co/datasets/HuggingFaceFW/fineweb/viewer/sample-10BT. We split 1% of the data as the corresponding development set. Models are trained for one epoch with a cosine scheduler applied and a 2000 step warm-up. Refer to Supplementary Material, Section 1.2 for details on how we process the training dataset.
@@
Besides reporting the training loss and perplexity results, we perform zero-shot evaluation on general language modeling tasks using the lm_eval benchmark (Gao et al., 2024). Specifically, we evaluate on WinoGrande (Sakaguchi et al., 2019), ARC (Clark et al., 2018), PIQA (Bisk et al., 2020) and SciQ (Welbl et al., 2017), which span a range of reasoning and question answering tasks across diverse domains.
Figure 2: Comparison of our DQT and other baselines across different model sizes and training datasets. The horizontal axis represents the training steps while the vertical axis represents the training loss. As model size increases, the performance of our DQT models, especially DQT 8 bit, become more comparable to and even better than the reproduced BitNet b1.58.
@@
We pre-train the 130M models on 4 NVIDIA A100 80GB GPUs, the 330M models on 8 NVIDIA GH200 Grace Hopper Superchips, and the 1B models on 16 GH200 Superchips. To remain hardware-agnostic, we simulate our quantization approach under FP32, BF16, and FP8 environments. We use MS-AMP [^4] https://github.com/Azure/MS-AMP for FP8 experiments with optimization level set to MS-AMP O2. While we recognize that FP8 is still not true ternary precision, it offers a practical compromise under current hardware constraints, as done in many prior works.
@@ 4.2 Main Results
We first present the training loss for DQT variants, our reproduced LLaMA (FP32) and BitNet b1.58 in Figure 2, including different model sizes and training datasets under the FP32 environment. In all subfigures of Figure 2, the blue line represents the ternary implementation of DQT, where weight matrices are always constrained to $\{-1,0,1\}$. The green line denotes the 8-bit (INT8) implementation of DQT (in the format of FP32 due to simulation). The orange line corresponds to the reproduced BitNet b1.58, which utilizes high-precision information during training. Finally, the red line indicates the standard FP32 LLaMA implementation.
We first examine the blue lines: although there remains a performance gap compared to higher-precision models, the ternary DQT implementation still demonstrates the ability to converge. Across all model sizes and training datasets, we observe that standard FP32 models (red lines) consistently achieve the best performance, as they are not subject to any quantization. BitNet (orange lines) consistently delivers performance close to FP32, benefiting from high-precision updates during training. Finally, for our DQT 8 bit implementation (green lines), the performance gap with BitNet b1.58 narrows as model size increases. Particularly, in the 1B models, DQT 8 bit models surpass BitNet b1.58, and this trend persists even when training on the larger FineWeb dataset. A clearer non-logarithmic comparison between the 1B DQT 8-bit model and BitNet b1.58 is provided in the Supplementary Material, Figure 1.
@@ 4.3 Low Memory Experiments
We additionally conduct two types of experiments to evaluate DQT’s performance given reduced GPU memory, simulating resource-limited environments. First, we assess performance in low-precision formats using BF16 and FP8  [^5] In our notation,“n-bit” refers to the quantization level of the model weights (INT-n), while the labels such as FP32, BF16, or FP8 refer to the precision of the environment. Note that any “n-bit” can be simulated in these precisions.. Second, we examine the effect of memory-efficient optimizers, as standard AdamW maintains two high-precision states (momentum and variance) for each parameter, contributing significantly to memory usage. To demonstrate DQT’s low reliance on high-precision information, we specifically choose Adafactor (Shazeer and Stern, 2018), which eliminates the need for fully storing these additional states. Note that we conduct experiments with Adafactor in BF16 and FP8 formats, as its benefits are most pronounced in low-precision environments.
Figure 3 presents the GPU memory usage and corresponding development set loss at the end of training for 130M and 1B models trained on the Wikipedia dataset. The y-axis indicates the percentage of GPU memory used on a single GH200 GPU under our experimental settings. We explain how we acquire the percentage in Supplementary Material, Section 1.4. As shown in the figure, BitNet models experience clear performance degradation as GPU memory usage decreases in BF16 and FP8 settings, for both 130M and 1B scales. This is expected, since BitNet still relies on high-precision information for effective weight updates. In contrast, our DQT 8-bit models maintain robust performance under reduced precision. For both 130M and 1B models, the performance drop is less than 0.1 in loss, demonstrating DQT’s resilience to lower GPU memory usage and its minimal dependence on high-precision information. This robustness persists even when applying memory-efficient optimizers such as Adafactor in both BF16 and FP8 settings. These results suggest that DQT models can be effectively trained even in memory-constrained environments, showing consistent performance across both low-precision formats and lightweight optimization methods. We believe that DQT can be further extended to even lower-precision settings while maintaining its effectiveness.
@@ 4.4 Impact of the Bit Width in DQT
Figure 4: Comparison of bit widths in DQT. Higher $n$-bit results in better performance.
Next, we proceed to examine the impact of varying the number of bits in DQT. Specifically, we experiment with $n$ in {1.58, 3, 4, 8} on 130M models trained with the Wikipedia dataset and 1B models trained with the FineWeb dataset under FP32. The corresponding results are presented in Figure 4. We select the optimal learning rate for each model.